# Spectral Properties of Irradiated Circumbinary Disks around Binary Black Holes Governed by Hydrogen Opacities Dependent on Temperature and Density

### v1

@@ 4.2 Circumbinary Disk Spectra
Figure 4 shows how the CBD spectrum changes when the outer disk radius $r_{\rm out}$ varies. Panels (a), (b), and (c) correspond to $r_{\rm out}=10^{4}\,r_{\rm in}$ (fiducial case), $10^{3}\,r_{\rm in}$, and $10^{5}\,r_{\rm in}$, respectively. The black, blue, green, and red curves represent Models 0, 1, 2, and 3, respectively. As $r_{\rm out}$ increases, the IR peak becomes stronger. Panel (c), which adopts the largest outer radius, shows the most prominent IR peak. This is because a larger outer radius allows the two minidisks to irradiate a wider region of the outer CBD. The radiation from the minidisks is absorbed and then re-emitted as IR light, increasing the total IR luminosity. In contrast, Panel (b), for the smallest outer radius, shows a much weaker IR peak due to its limited irradiated area.
The spectral differences between the models also become more noticeable in the IR region as $r_{\rm out}$ increases. This is because the opacities adopted in each model mainly affect the efficiency of reprocessing the irradiating photons. When irradiation covers a larger portion of the disk, these opacity effects become more significant and lead to clearer differences among the models. Therefore, both the size of the CBD and the strength of irradiation are key to understanding how different types of opacity—such as electron scattering, free–free absorption, and bound–free absorption—influence the emergent spectrum.
4.3 Triple disk spectra
We compute the SED of a triple disk system consisting of the CPD, the CSD, and the surrounding CBD. The spectral luminosity of each minidisk is calculated using the relation $L_{\nu,i}=4\pi D^{2}S_{\nu,i}$, where $S_{\nu,i}$ is the flux density emitted from the CPD ($i=1$) or the CSD ($i=2$). This flux density is computed using the same integrand as in equation (53), but it is integrated over a distinct radial range, from the ISCO radius, $r_{{\rm in},i}=6GM_{i}/c^{2}$, to the outer edge of each disk. The outer edge is defined as a fixed fraction (typically $\sim 0.5$) of the Hill radius, which represents the tidal truncation limit for each minidisk. This choice reflects the results of numerical simulations showing that tidal truncation limits the disk extent well within the Hill sphere (e.g., Martin & Lubow 2011; Paschalidis et al. 2021). The temperature profile of each minidisk follows the standard thin-disk scaling, $T_{i}(r)\propto r^{-3/4}$, where $r_{{\rm in},i}\leq{r}\leq{r_{{\rm out},i}}$.
@@ 4.3 Triple disk spectra
4.4 Observational implications
@@
Figure 6: Comparison of Subaru, JWST, and Swift/XRT sensitivity limits with theoretical triple disk spectra at different source distances. The vertical axis shows the flux density (in erg s-1 cm-2 Hz-1), and the horizontal axis represents the frequency (in Hz); both are plotted on logarithmic scales. All models adopt fiducial parameters of $M=100\,M_{\odot}$ and $r_{\rm out}=10^{4}\,r_{\rm in}$. The solid black line denotes Model 0, while the solid red and green lines represent Model 3 at distances of $D=10$ Mpc and $D=100$ Mpc, respectively. The dashed horizontal curves indicate the flux sensitivity limits of JWST/NIRCam filters (F [^4] W through F070W), Subaru/MOIRCS and HSC filters ($Y$, $z$, $i2$, $r2$, and $g$), and Swift/XRT in the 0.2–10 keV X-ray band. This comparison highlights the observability of CBD features in the IR-optical-X-ray bands and demonstrates that, for $D=100$ Mpc, only the X-ray emission (primarily from the CPD and CSD) is detectable, whereas at $D=10$ Mpc, the irradiated CBD in Model 3 becomes marginally detectable in the Subaru/HSC $r2$ band.
@@
Figure 7: Rosseland mean opacities at a fixed density of $\rho=10^{-10}\,\mathrm{g\,cm^{-3}}$. The dashed green, magenta, and orange lines represent the contributions from free-free absorption, bound-free absorption, and electron scattering for hydrogen gas, respectively. The solid black line represents the total analytical opacity, calculated as $\kappa_{\rm ff}+\kappa_{\rm bf}+x_{e}\,\kappa_{\rm es}$, where the ionization fraction $x_{e}$ is determined self-consistently from the Saha equation. The electron scattering term $\kappa_{\rm es}$ is multiplied by $x_{e}$ to account for its dependence on free electrons. The solid blue line indicates the tabulated opacity from the Opacity Project (Seaton, 2005), an international collaboration to calculate the extensive atomic data required to estimate stellar envelope opacities, and from (Semenov et al., 2003), including dust opacity based on normal silicate grains under the assumption of homogeneous spheres.
5 Discussion
@@
This section discusses two key caveats of our models. First, we examine how realistic opacity prescriptions affect the emergent spectrum from an irradiated CBD, focusing on discrepancies between our models and more detailed opacities available in tabulated forms. Second, we assess the detectability of GWs emitted from such systems by current and future GW observatories, based on the binary orbital parameters.
@@
In Figure 7, we compare the Rosseland mean opacity computed from our models—including free-free absorption, bound-free absorption, and electron scattering processes—with the tabulated values from the Opacity Project (Seaton, 2005), which provides extensive atomic data for stellar envelopes, and from Semenov et al. (2003), widely used in the context of protoplanetary disk modeling. Our model employs cross-section formulae and ionization fractions derived from the Saha equation, assuming pure hydrogen gas at a fixed density of $\rho=10^{-10}~{}{\rm g\,cm^{-3}}$. The comparison reveals significant discrepancies, particularly in the temperature range of $T\sim 10^{3}$$10^{4}$ K, where the tabulated Rosseland opacities are several orders of magnitude lower than those predicted by our analytical prescription that includes the bound-free absorption. This divergence arises because the tabulated opacities incorporate more realistic astrophysical conditions, such as grain sublimation, grain growth, and composition-dependent absorption features, all of which are absent from our simplified hydrogen-only framework.
@@
These differences in opacity have direct consequences for the emergent spectrum from irradiated CBDs. In the temperature range of $T\sim 10^{3}$$10^{4}$ K, the reduced opacity in the tabulated models is expected to enhance radiative cooling, allowing energy to be transported more efficiently to the disk surface and resulting in a higher surface temperature. Consequently, the overall spectral intensity, particularly in the far-IR and sub-millimeter bands dominated by reprocessed irradiation, increases.
@@
As a future direction, it will be essential to incorporate opacity tables (e.g., Semenov et al. 2003; Seaton 2005) directly into irradiated disk models, enabling systematic computation and comparison of CBD spectra under more realistic astrophysical conditions. This will provide a more robust framework for interpreting multi-wavelength observations and constraining the physical parameters of circumbinary environments.
If the binary system is undergoing orbital decay due to GW emission, the emitted GW signal is characterized by its frequency and strain amplitude, both determined by the binary’s orbital parameters. In our fiducial model, the GW frequency and strain are given by (Moore et al., 2015)
$f_{\rm GW}$ $=$ $\frac{\Omega}{\pi}\sim 7.2\times 10^{-3}\,{\rm Hz}\left(\frac{M}{100\,M_{\odot}}\right)^{-1},$ (54)
$h_{\rm c}$ $=$ $\frac{2(G\mathcal{M})^{5/3}(\pi f_{\rm GW})^{2/3}}{c^{4}D}\sim 1.2\times 10^{-22}\left(\frac{M}{100\,M_{\odot}}\right)\left(\frac{D}{10\,{\rm Mpc}}\right)^{-1},$ (55)
respectively, where $\mathcal{M}=(M_{1}M_{2})^{3/5}/(M_{1}+M_{2})^{1/5}$ is the chirp mass, and $D$ is the distance to the source.

### v2

@@ IV.2 Circumbinary Disk Spectra
Figure 5 shows how the CBD spectrum changes when the outer disk radius $r_{\rm out}$ varies. In this figure, panels (a), (b), and (c) correspond to $r_{\rm out}=10^{4}\,r_{\rm in}$ (fiducial case), $10^{3}\,r_{\rm in}$, and $10^{5}r_{\rm in}$, respectively. The black, blue, green, and red curves represent Models 0, 1, 2, and 3, respectively. As $r_{\rm out}$ increases, the irradiated area of the outer CBD becomes larger, the reprocessed radiation from the two minidisks is absorbed and re-emitted over a wider region, and the IR peak becomes stronger; panel (c), which has the largest outer radius, therefore shows the most prominent IR peak, whereas panel (a), with the smallest radius, shows the weakest IR excess. The spectral differences between the models also become more noticeable in the IR region as $r_{\rm out}$ increases, because the opacities adopted in each model mainly control how efficiently the irradiating photons are reprocessed. When irradiation covers a larger fraction of the disk, these opacity effects are amplified and lead to clearer separations among the model spectra.
These trends are also consistent with the scaling of the irradiation heating rate derived in Section III.4. In the $\tau_{3}$-opacity regime that we use to describe the outer CBD, equation (57) shows that the local irradiation heating rate evaluated at the outer boundary decreases as $Q_{\rm irr}(\xi_{\rm out})\propto\xi_{\rm out}^{-69/38}$ when the other parameters are kept fixed. Since a larger outer radius increases the reprocessing area and includes cooler annuli that dominate the lowest-frequency emission, $\xi_{\rm out}$ mainly sets how far the IR bump extends to low frequencies and, through the radial integration, how strong it appears the overall irradiation-powered IR bump.
Finally, we examine how the CBD spectrum depends on the binary semi-major axis at a fixed physical outer radius. As equation (3) shows, the inner edge of the CBD scales with the semi-major axis as $r_{\rm in}=C_{\rm gap}a=C_{\rm gap}a_{0}r_{\rm S}$, so that the outer radius is $r_{\rm out}=\xi_{\rm out}r_{\rm in}\propto\xi_{\rm out}a_{0}$ for fixed $C_{\rm gap}$ and $M$. Using the fiducial model with $a_{0}=1000$ and $\xi_{\rm out}=10^{4}$ as a reference, we therefore adopt $\xi_{\rm out}=10^{5}$ for a more compact binary with $a_{0}=100$, so that $r_{\rm out}$ is identical in both cases. This choice allows us to isolate the effect of changing $a_{0}$ without introducing any change in the emitting area of the outer disk.
The dynamical coupling between the CBD and the binary is governed by whether the viscous inflow speed at the cavity rim can keep up with the GW-driven inspiral rate (see Armitage & Natarajan 2002; Dittmann et al. 2023). To verify that the binary separations explored here correspond to dynamically disk-coupled configurations, we estimate the decoupling radius by equating the GW-driven inspiral speed to the viscous inflow speed at $r_{\rm in}$ as
$a_{\rm dec}$ $=$ $\frac{1}{2}\left[\frac{128}{15}\,\frac{q}{(1+q)^{2}}\,\frac{\sqrt{C_{\rm gap}}}{\alpha_{\rm SS}\,Y_{\rm in}^{2}}\right]^{2/5}r_{\rm S}$
$\sim$ $70\,r_{\rm S}\,\left(\frac{\mathcal{Q}}{1/4}\right)^{2/5}\left(\frac{C_{\rm gap}}{2}\right)^{1/5}\left(\frac{\alpha_{\rm SS}}{0.1}\right)^{-2/5}\left(\frac{Y_{\rm in}}{0.01}\right)^{-4/5}$
(e.g., equation (3) of D’Orazio et al. 2025), where $Y_{\rm in}=(H/r)_{r=r_{\rm in}}$ and $\mathcal{Q}\equiv q/(1+q)^{2}$. For our fiducial parameters, both $a_{0}=100$ and $a_{0}=1000$ exceed $a_{\rm dec}/r_{\rm S}$ (with $a_{0}=100$ being closer to the transition). Therefore, the binaries considered in this paper remain dynamically coupled to the CBD.
Figure 6 compares the CBD spectra for Models 0 and 3, while Models 1 and 2 always lie between these two cases and are omitted for clarity. For both opacity prescriptions, decreasing $a_{0}$ moves the high-frequency peak to a higher frequency and increases its amplitude because the inner edge ($r_{\rm in}$) becomes smaller and hotter. This behavior is consistent with equation (56), in which the viscous heating rate $Q_{\rm vis}$ near $r_{\rm in}$ scales as $(C_{\rm gap}a_{0})^{-3}$.
As noted above, we choose $(a_{0},\xi_{\rm out})=(1000,10^{4})$ and $(100,10^{5})$ so that the physical outer radius $r_{\rm out}=\xi_{\rm out}r_{\rm in}$ is the same in both cases. Consequently, the irradiated area of the outer CBD and the characteristic incident flux $(L_{1}+L_{2})/(4\pi r_{\rm out}^{2})$ are similar for the two models, as is the total energy available for reprocessing in the CBD.
Equation (57) describes how the local irradiation heating rate $Q_{\rm irr}$ scales at large radii when $\xi_{\rm out}$ is varied at fixed model parameters. In our setup, however, $a_{0}$ and $\xi_{\rm out}$ are varied together so that $r_{\rm out}$ stays fixed, cancelling much of the formal dependence of $Q_{\rm irr}$ on $a_{0}$ and $\xi_{\rm out}$. The surface temperature responds only as $T_{\rm eff}\propto Q_{\rm irr}^{1/4}$, which further weakens the remaining dependence. In addition, the IR bump is produced by emission integrated over a broad range of radii rather than only at $r_{\rm out}$. This means that enhancements of $Q_{\rm irr}$ near the outer edge are diluted in the total spectrum. For these reasons, the height of the irradiation-powered IR bump varies only weakly between the two values of $a_{0}$. The small remaining differences in the IR range appear mainly in Model 3, where bound-free opacity slightly increases the reprocessing efficiency in the cool outer layers, producing a somewhat broader low-frequency shoulder. In this comparison, changing $a_{0}$ primarily affects the inner, viscously heated component at fixed $r_{\rm out}$. The structure of the IR bump is governed by the disk size and the opacity model rather than by $a_{0}$ itself.
Figure 6: CBD spectra for different binary separations at a fixed physical outer radius. The normalized luminosity $\nu L_{\nu}/L_{\rm Edd}$ is plotted as a function of frequency $\nu$. The four curves show Models 0 and 3, which are described in Table 1, for two combinations of the dimensionless parameters: $(a_{0},\xi_{\rm out})=(1000,10^{4})$ and $(100,10^{5})$. The values of $\xi_{\rm out}$ are chosen so that the physical outer radius $r_{\rm out}=\xi_{\rm out}r_{\rm in}\propto\xi_{\rm out}a_{0}$ is the same in both cases.
IV.3 Triple disk spectra
We compute the SED of a triple disk system consisting of the CPD and CSD (hereafter “minidisks”), and the surrounding CBD. The minidisks’ SED are modeled as surface-emission integrals over local blackbodies. In our framework, the emergent multi-color blackbody spectrum from each minidisk is controlled by the surface temperature profile of each minidisk, which is determined to be proportional to $r^{-3/4}$ for given black hole mass and mass accretion rate by the local dissipation rate of the standard disk model (Shakura & Sunyaev, 1973; Pringle, 1981). The pressure (gas- vs. radiation-pressure dominated) and opacity regimes modify primarily the vertical structure and midplane temperature but do not alter the resulting multi-color blackbody spectrum (e.g., Kato et al. 2008).
The spectral luminosity of each minidisk is calculated using the relation $L_{\nu,i}=4\pi D_{\rm L}^{2}S_{\nu,i}$, where $S_{\nu,i}$ is the flux density emitted from the CPD ($i=1$) or the CSD ($i=2$). This flux density is computed using the same integrand as in equation (58), but it is integrated over a distinct radial range, from the ISCO radius, $r_{{\rm in},i}=6GM_{i}/c^{2}$, to the outer edge of each disk. The outer edge radius is defined as a fixed fraction (typically $\sim 0.5$) of the Hill radius, which represents the tidal truncation limit for an accretion disk around a primary object in a binary system (e.g., Martin & Lubow 2011; Paschalidis et al. 2021). For a BBH with minidisks, their tidal truncation radii are numerically calibrated to be $r_{\rm out,\it i}\simeq 0.3-0.4\,a$ (e.g., Farris et al. 2014; Miranda et al. 2017; Duffell et al. 2020; Muñoz et al. 2020; D’Orazio & Duffell 2021). At the minidisk photosphere, we approximate the opacity as electron-scattering dominated, which is appropriate for the hot minidisk temperatures ($\gtrsim 10^{5}\,{\rm K}$) in our parameter range. Electron scattering can harden the emergent spectrum relative to a blackbody (Shimura & Takahara, 1995). For simplicity, we neglect this effect and model each minidisk as an uncorrected multi-color blackbody.
The triple disk spectrum is then given by the sum of CBD, CPD, and CSD specific luminosities as
$L_{\nu}=L_{\nu,{\rm CBD}}+L_{\nu,1}+L_{\nu,2}$
where $L_{\nu,{\rm CBD}}$ is obtained from the CBD vertical solution including irradiation and our opacity prescriptions described in Sections IV.1 and IV.2. In practice, the composite SED shows a high–frequency peak from the hot inner minidisks and an IR–optical excess shaped by CBD irradiation/reprocessing.
@@ IV.3 Triple disk spectra
Figure 8: Comparison of Subaru, JWST, and Swift/XRT sensitivity limits with theoretical triple disk spectra at different source distances. The vertical axis shows the flux density (in erg s-1 cm-2 Hz-1), and the horizontal axis represents the frequency (in Hz); both are plotted on logarithmic scales. All opacity models adopt fiducial parameters of $M=100\,M_{\odot}$ and $r_{\rm out}=10^{4}\,r_{\rm in}$. The solid black line denotes Model 0, while the solid red and green lines represent Model 3 at distances of $D_{\rm L}=10\,\mathrm{Mpc}$ and $D_{\rm L}=100\,\mathrm{Mpc}$, respectively. The dashed horizontal curves indicate the flux sensitivity limits of JWST/NIRCam filters (F [^4] W through F070W), Subaru/MOIRCS and HSC filters ($Y$, $z$, $i2$, $r2$, and $g$), and Swift/XRT in the 0.2–10 keV X-ray band. This comparison highlights the observability of CBD features in the IR-optical-X-ray bands and demonstrates that, for $D_{\rm L}=100\,\mathrm{Mpc}$, only the X-ray emission (primarily from the CPD and CSD) is detectable, whereas at $D_{\rm L}=10\,\mathrm{Mpc}$, the irradiated CBD in Model 3 becomes marginally detectable in the Subaru/HSC $r2$ band.
IV.4 Observational implications
@@
IV.5 Gravitational wave detectability
Since the binary loses orbital energy and angular momentum through gravitational radiation, the waveform is determined by the orbital parameters. In the quasi-circular regime the GW frequency increases slowly while the strain grows as the orbit shrinks (Sathyaprakash & Schutz, 2009). For our purpose we describe this slow chirp with a frequency-domain treatment that connects the orbital evolution to the observed characteristic strain and the mission-averaged signal-to-noise ratio (SNR).
The characteristic strain under the stationary phase approximation (SPA) is given by (e.g., Sathyaprakash & Schutz 2009; Moore et al. 2015)
$h_{\rm c}(f)=\frac{8}{\sqrt{5}}T_{\rm obs}^{1/2}\frac{\bigl(G\mathcal{M}\bigr)^{5/3}}{D_{\rm L}c^{4}}\pi^{2/3}f^{7/6},$ (60)
where $\mathcal{M}=\bigr[q/(1+q)^{2}\bigr]^{3/5}M$ is the chirp mass and the GW frequency is given by
$f=\frac{2}{P_{\rm orb}}=\frac{1}{\pi}\sqrt{\frac{GM}{a^{3}}}.$ (61)
The frequency evolution in the inspiral is then obtained by
$\frac{df}{dt}=\frac{96}{5}\,\pi^{8/3}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{5/3}f^{11/3},$ (62)
which we use to determine the frequency interval swept during the observing time $T_{\rm obs}$.
We compute the mission-averaged SNR using the total LISA noise $S_{\mathrm{tot}}(f)$ of Robson et al. (2019), which combines the instrument and Galactic confusion contributions:
$\bar{\rho}_{\rm SNR}^{2}=\int_{f_{0}}^{f_{0}+\Delta{f}}\frac{h_{\rm c}^{2}(f)}{fS_{\mathrm{tot}}(f)}\,\mathrm{d}f,$ (63)
where $f_{0}$ is the entry frequency at the chosen orbital separation $a=a_{0}r_{\rm S}$, and $\Delta f\simeq(df/dt)\,T_{\rm obs}$ if the chirp is slow. To complement the EM predictions, we evaluate the LISA detectability of three representative mass cases: C1: $M=10\,M_{\odot}$, C2: $M=100\,M_{\odot}$, and C3: $M=1000\,M_{\odot}$. Throughout this subsection C1–C3 refer to the mass cases and should not be confused with the opacity models 1–3 defined earlier. We fix the semi-major axis at $a=1000\,r_{\rm S}$ and the luminosity distance at $D_{\rm L}=10~\mathrm{Mpc}$ for all three cases. As estimated at the end of Section IV.2, the decoupling radius is typically $a_{\rm dec}\sim 70\,r_{\rm S}$ for our fiducial parameters. The $a_{0}=1000$ cases, therefore, remain in the pre-decoupling regime. The resulting quantities used for the three masses are listed in Table 2, and their locations relative to the LISA instrument, confusion, and total amplitude spectral densities are shown in Figure 9.
It is noted from the figure that C2 and C3 are detected with high confidence ($\bar{\rho}_{\rm SNR}\simeq 1.98\times 10^{2}$ and $4.72\times 10^{1}$) under the conventional detection threshold $\bar{\rho}_{\rm SNR}\gtrsim 8$, whereas C1 remains sub-threshold ($\bar{\rho}_{\rm SNR}\simeq 3.64$). For the C2 case, the binary enters the LISA band at a GW frequency of $f_{0}\simeq 7.2~\mathrm{mHz}$, which lies very close to the frequency where the total noise amplitude $S_{\rm tot}^{1/2}(f)$ reaches its minimum. During the mission, the signal then sweeps across a logarithmic frequency interval of $\Delta\ln f\simeq 7.7\times 10^{-2}$, so it spends a significant time in this most sensitive part of the band. This combination of near–optimal frequency and non-negligible bandwidth yields the largest SNR among the three cases. C1 lies on the high-frequency rise where $S_{\rm inst}^{1/2}$ increases, whereas C3 lies in the sub-mHz region dominated by the confusion foreground; both suffer reduced SNR relative to C2.
Since $h_{c}\propto D_{\rm L}^{-1}$, the SNR scales as $\bar{\rho}_{\rm SNR}\propto D_{\rm L}^{-1}$, at $D_{\rm L}=100$ Mpc the three cases give roughly $\bar{\rho}_{\rm SNR}\approx 0.36$ (C1), $19.8$ (C2), and $4.7$ (C3). For fixed $a/r_{\rm S}$, the available $\Delta\ln f$ decreases approximately with increasing $M$ over a finite observing time, which partly offsets the intrinsic rise of $h_{c}$ with mass.
In our calculation of $\bar{\rho}_{\rm SNR}$ we use the SPA and integrate the signal power along the entire inspiral frequency track. For the noise, we adopt the total LISA power spectral density $S_{\rm tot}(f)$, which includes both the instrumental noise and the Galactic confusion noise. Even if we vary the assumed observing time within a plausible range of a few years, or modify the level of confusion–noise subtraction within realistic limits, the resulting SNRs change only by factors of order unity (typically less than a factor of a few). The qualitative ranking of the three cases, $C2>C3\gg C1$, remains unchanged.
Among the three cases, C2 is the most promising for joint GW–EM observations: it is the loudest in LISA and, in our irradiated-CBD framework, it exhibits a strong IR/optical reprocessing signature. For very nearby events, C1 becomes detectable only at distances of a few Mpc, while C3 remains detectable within $\lesssim 10\,\mathrm{Mpc}$, enabling direct tests of the opacity-dependent CBD spectra presented in Sections IV.2IV.4
Case Binary mass ($M$) GW frequency ($f$) $\Delta\ln f$ Characteristic strain ($h_{\rm c}$) SNR ($\mathcal{\bar{\rho}}_{\rm SNR}$)
$M_{\odot}$ $\rm Hz$} - - -
C1 $10$ $2\times 10^{-2} 7.7$$7\times 10^{-1}$ $6.5\times 10^{-20}$ $3.6$
C2 $100$ $2\times 10^{-3} 7.7$$7\times 10^{-2}$ $2.0\times 10^{-19}$ $198$
C3 $1000$ $2\times 10^{-4} 7.7$$7\times 10^{-3}$ $6.5\times 10^{-19}$ $47$
Table 2: GW observables and SNR for the three BBHs with different masses, evaluated at a common orbital separation $a_{0}=1000$. Columns list the model label, binary mass $M$, GW frequency $f$, characteristic strain $h_{c}$ (quadrupole estimate at a common luminosity distance $D_{\rm L}=10\,\mathrm{Mpc}$, and SNR $\bar{\rho}_{\rm SNR}$ for LISA. All binaries are assumed circular, face-on, and equal-mass.
Figure 9: Comparison of the LISA sensitivity $S_{\rm tot}^{1/2}(f)$ with the gravitational-wave signals for three different masses: C1 (black), C2 (red), and C3 (blue). The horizontal axis is the GW frequency $f$ in units of Hz, and the vertical axis is the characteristic strain $h_{\rm c}$. Colored points mark $h_{\rm c}$ evaluated at $q=1$, $a=1000\,r_{\rm S}$, and $D_{\rm L}=10\,\mathrm{Mpc}$, whereas the black line shows $\sqrt{f\,S_{\rm tot}(f)}$. All values assume circular, face-on, equal-mass binaries. It is noted that $S_{\rm tot}(f)$ is the sum of the amplitude spectral densities of the LISA instrument noise $S_{\rm inst}(f)$ and Galactic confusion noise $S_{\rm c}(f)$.
V Discussion
@@
Figure 10: Comparison of analytic hydrogen opacities and tabulated Rosseland-mean opacities at a fixed density of $\rho=10^{-10}\,{\rm g\,cm^{-3}}$. The dashed black, red, and blue lines represent the contributions from electron scattering, free-free absorption, and bound-free absorption for hydrogen gas, respectively. The solid black line represents the total analytical opacity, calculated as $x_{e}\kappa_{\rm es}+\kappa_{\rm ff}+\kappa_{\rm bf}$, where the ionization fraction $x_{e}$ is determined self-consistently from the Saha equation. The solid green line indicates the tabulated opacity from the Opacity Project (Seaton, 2005), an international collaboration to calculate the extensive atomic data required to estimate stellar envelope opacities, and from (Semenov et al., 2003), including dust opacity based on normal silicate grains under the assumption of homogeneous spheres.
@@
This section discusses several key caveats of our models. First, we examine how realistic opacity prescriptions affect the emergent spectrum from an irradiated CBD, focusing on discrepancies between our opacity models and more detailed opacities available in tabulated forms. Second, we assess the detectability of GWs emitted from such systems by current and future GW observatories, based on the binary orbital parameters.
V.1 Opacity Prescriptions and Comparison with Tabulated Opacities
@@
We now compare our analytic hydrogen opacities with widely used tabulated opacity tables. Figure 10 compares the Rosseland mean opacity computed from our opacity models, which include free–free absorption, bound–free absorption, and electron scattering processes, with the tabulated values from the Opacity Project (Seaton, 2005), which provides extensive atomic data for stellar envelopes, and from Semenov et al. (2003), which are widely used in protoplanetary disk modeling. In addition to tabulated opacities, protostellar disk studies commonly use piecewise, temperature-dependent fits for the low-temperature Rosseland-mean opacity (e.g., Bell & Lin 1994; Bell et al. 1997). Our model employs cross-section formulas and ionization fractions derived from the Saha equation, assuming a pure hydrogen composition at a fixed density of $\rho=10^{-10}~{\rm g\,cm^{-3}}$. The comparison reveals significant discrepancies, particularly in the temperature range $T\sim 10^{3}$$10^{4}~{\rm K}$. In this range, the tabulated Rosseland opacities are several orders of magnitude lower than those predicted by our analytical prescription that includes bound–free absorption. This difference arises because the tabulated opacities incorporate more realistic astrophysical conditions, such as grain sublimation, grain growth, and composition-dependent absorption features, all of which are absent from our simplified hydrogen-only framework.
@@
These differences in opacity have direct consequences for the emergent spectrum from irradiated CBDs. In the temperature range of $T\sim 10^{3}$$10^{4}$ K, the reduced opacity in the tabulated models is expected to enhance radiative cooling, so that energy is transported more efficiently to the disk surface and the surface temperature becomes higher. Consequently, the overall spectral intensity, particularly in the far-IR and submillimeter bands dominated by reprocessed irradiation, increases.
In the supermassive regime, equation (57) implies that the irradiation–dominated heating rate scales as $Q_{\rm irr}\propto M^{-21/19}$. Combined with $Q_{\rm irr}=2\sigma T_{\rm s}^{4}$, this yields $T_{\rm s}\propto M^{-21/76}$. Accordingly, extending our fiducial $100\,M_{\odot}$ model to $M\sim 10^{6}$$10^{8}\,M_{\odot}$ would reduce the outer surface temperature by more than an order of magnitude, from $\sim 10^{3}$$10^{4}$ K to $T_{\rm s}\lesssim$ a few $\times 10^{2}$ K. At such temperatures, dust and metal opacities dominate over hydrogen bound–free and free–free opacity, and full tabulated Rosseland means (e.g. Semenov et al. 2003; Ferguson et al. 2005; Seaton 2005) are required for reliable spectra. A self-consistent treatment of the circumbinary SEDs of SMBH binaries therefore demands replacing the analytic Kramers prescriptions employed here with composition-dependent opacity tables, and we defer this extension to future work, restricting the present quantitative SED calculations to stellar- and intermediate-mass BBHs while summarizing only the SMBH mass-scaling.
@@
As a future direction, it will be essential to incorporate opacity tables (e.g., Semenov et al. 2003; Seaton 2005) directly into irradiated disk models, enabling systematic computation and comparison of CBD spectra under more realistic astrophysical conditions. This will provide a more robust framework for interpreting multiwavelength observations and constraining the physical parameters of circumbinary environments.
V.2 Orbital Decay, Circumbinary Disk Decoupling, and GW Detection Across Mass Scales